An airplane took 2.5 hours to fly 625 miles with the wind. It took 4 hours and 10 minutes to make the return trip, against the same wind. Find the wind speed and the speed of the plane in still air.
Against a head wind, Jeff computes his flight time for a trip of 2,900 miles at 5 hours. The flight would take 4 hours and 50 minutes if the head wind was half as much. Find the head wind and the planes air speed.
In 1927 Charles Lindburg flew the Spirit of St. Louis from NYC to Paris at an average speed of 107.4mi/h. 11 years later, Howard Hughes flew the same route, averaged 217.1 mil/h, and took 16 hours and 57 minutes less time. Find the length of their route.

Answer by piglet162431(90)

let statements:
let x be the be speed of the plane in still air.
Let y be the speed of the wind.
since d=rt, we can come upwith a system of equations.
625=2.5(x+y)
625=4 1/6(x-y)
The first equation could be simplied into 250=x+y.
the second equation, multiply both sides to the reciprocal of 4 1/6, which is 6/25:
150=x-y
now, line them up, and use addition method.
x-y=150
x+y=250
2x=400
x=200
Plug this value back for y. 200-y=150, y=50.
The speed of the plane in still is 200 m/h, whhile the speed of the wind is 50 m/h.